OpenStudy (anonymous):

What is the equation of a line that is parallel to the line y=3/4x+2 and passes through the point (-2,8) ?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

1.) y= 3/4x + 9/2 2.) y= 3/4x - 19/2 3.) y= 3/4 + 19/2 4.) y= 3/4x + 19/2

OpenStudy (skullpatrol):